# Integrated Vision and Language Lab, KAIST, South Korea

### v1

@@ 5 Conclusion
This paper proposed a multilingual VSR method using a single model. Specifically, we proposed to use visual speech unit modeling to mitigate the huge computational loads in building massive multilingual visual speech processing models. To this end, we introduced multilingual AV-HuBERT (mAV-HuBERT) trained on audio-visual data of nine languages. With the visual speech units, we can greatly reduce the data size and effectively pre-train the VSR model on large-scale multilingual VSR databases. By analyzing the visual speech unit, we validated it contains linguistic information and enables visual speech modeling using discrete inputs. To complement visual speech information with audio, we proposed curriculum learning by gradually increasing the task difficulty. By finetuning the model on continuous features, we set new state-of-the-art multilingual VSR performances by achieving comparable VSR performances with the previous language-specific VSR models.
Appendix A Appendix
A.1 Visualization of Speech Units
@@
A.2 Dataset Statistics for Each Language
The dataset statistics for each language are shown in Table 8. Please note that there are only 181,034 non-English human-labeled videos (i.e., mTEDx). Therefore, we increase the quantity of labeled data by utilizing the automatic labels proposed by [20, 39]. With this, we can construct 2,014,212 multilingual video-text paired data. The number of these automatic labels can be found in the ‘Auto-labeled # of Video’ column in the table.
Table 8: Statistics of datasets used in this work including automatic labels.
@@ A.2 Dataset Statistics for Each Language
A.3 Detailed Training Setup
@@
Table 9: Details of hyperparameters used in training.
  Pre-training (mAV-HuBERT) Pre-training (Visual speech unit to text) Fine-tuning (Multilingual)
# of epochs 40 11 8
# of steps 350,000 60,000 120,000
# of frozen steps - - 10,000
# of GPUs 64 32 32
Max frames / batch 1000 3000 1000
LR scheduler polynomial decay tri-stage tri-stage
warmup updates 48,000 15,000 15,000
peak learning rate 2e-3 1e-3 4e-4
Adam $(\beta_{1},\beta_{2})$ (0.9, 0.98) (0.9, 0.98) (0.9, 0.98)
A.4 Detailed Speaker Verification Setup
We utilize a pre-trained speaker recognition model of [81] and train the model with different inputs. To match the input size with the model, we use one additional embedding layer for both audio speech units and visual speech units. When we use the speech features from mAV-HuBERT, we utilize one additional linear layer for both audio speech features and visual speech features. For training, we utilize data less than 20 seconds in VoxCeleb2 [68] dev set. The speaker verification is performed on the test data of VoxCeleb2 having less than 20 seconds, where each test sample is assigned one positive sample and one negative sample. Therefore, 27,816 positive pairs and negative pairs are utilized respectively, thus random prediction yields 50% EER.
A.5 Examples of Predicted Sentences
@@
Fig. 5: Visualization of speech units. Each boundary represents a single unit and the same color represents the same phoneme or phoneme family. (a) Audio speech unit. (b) Visual speech unit.
Fig. 6: Example sentences predicted from the single proposed multilingual VSR model on LRS3 and mTEDx test set. The Red and Blue indicate deletion and wrong predicted words, respectively.

### v2

@@ 5. Conclusion
In this paper, we proposed an efficient multilingual VSR method using a single model. Specifically, we proposed to use visual speech units for pre-training a VSR model to mitigate the huge computational loads in building massive multilingual visual speech modeling. With the proposed strategy, we can greatly reduce the data size and effectively pre-train the VSR model on large-scale multilingual VSR databases. By analyzing the visual speech unit, we validated it contains linguistic information and enables visual speech modeling using discrete inputs. To complement visual speech information with audio, we proposed curriculum learning by gradually increasing the task difficulty. Finally, by finetuning the model on continuous features, we set new state-of-the-art multilingual VSR performances by achieving comparable VSR performances with the previous language-specific VSR models. To the best of our knowledge, this is the first work exploring the multilingual VSR and employing visual speech units as inputs of VSR systems.
Appendix A Visualization of Speech Units
@@
Figure 5. Visualization of speech units. Each boundary represents a single unit and the same color represents the same phoneme or phoneme family. (a) Audio speech unit. (b) Visual speech unit.
Appendix B Dataset Statistics for Each Language
The dataset statistics for each language are shown in Table 8. Please note that there are only 181,034 non-English human-labeled videos (i.e., mTEDx). Therefore, we increase the quantity of labeled data by utilizing the automatic labels proposed by (Yeo et al., 2023b; Ma et al., 2023). With this, we can construct 2,014,212 multilingual video-text paired data. The number of these automatic labels can be found in the ‘Auto-labeled # of Video’ column in the table.
@@ Appendix B Dataset Statistics for Each Language
Table 8. Data statistics of each language used in this work including automatic labels.
Appendix C Analysis of Linguistic Information in Visual Speech Units
To evaluate the extent to which linguistic information is maintained in visual speech units compared to continuous representations, we assess the speech recognition performances (WER) using five different input representations: raw audio, audio features, visual features, audio speech units, and visual speech units, on the LRS3 dataset. The results are shown in Table 9. In particular, we can confirm that discretization does not significantly compromise the linguistic content of the visual speech, with WERs ranging from 38.1% to 41.9%. Similarly, the audio speech units also demonstrate sufficient performance in retaining the linguistic content. By comparing these results with the speaker verification experiments in Table 4, we can confirm that the visual speech units contain linguistic information more than the speaker information.
2.7% 4.2% 38.1% 11.5% 41.9%
Table 9. WER comparisons using different inputs.
Method En Es It Fr Pt
AV $\rightarrow$ V 39.4% 69.5% 72.9% 78.0% 77.0%
Curriculum 37.5% 68.1% 70.3% 76.3% 74.0%
Table 10. Performance of different pre-training strategies.
Appendix D Effectiveness of Curriculum Learning
Instead of using curriculum learning, which starts the training with audio-visual speech units and gradually masks out the audio stream part, another training strategy could be transfer learning. This consists of two stages: pre-training the model with audio-visual speech units and then transferring to visual-only speech units. Both strategies share the objective of initially training the model with rich audio-visual speech information and then transforming the model to work with visual-only speech. To evaluate the effectiveness of these two strategies, we report the performances of 1) Transfer Learning, which involves the two-stage learning process, and 2) Curriculum Learning, which uses the masking strategy. The results are shown in Table 10. The performances of both methods are measured without finetuning on the continuous representations. By comparing these results, we confirm that the curriculum learning strategy showed relative improvements over the sequential training approach (i.e., transfer learning).
Appendix E Detailed Training Setup
@@
  Pre-training (mAV-HuBERT) Pre-training (Visual speech unit to text translation) Fine-tuning (Multilingual) Fine-tuning (Monolingual)
# of epochs 40 11 8 -
# of steps 350,000 60,000 120,000 60,000
# of frozen steps - - 10,000 7,200
# of GPUs 64 32 32 8
Max frames / batch 1000 3000 1000 1000
LR scheduler polynomial decay tri-stage tri-stage tri-stage
warmup updates 48,000 15,000 15,000 15,000
peak learning rate 2e-3 1e-3 4e-4 4e-4
Adam $(\beta_{1},\beta_{2})$ (0.9, 0.98) (0.9, 0.98) (0.9, 0.98) (0.9, 0.98)
Table 11. Details of hyperparameters used in training.
Appendix F Examples of Predicted Sentences
@@
Figure 6. Example sentences predicted from the single proposed multilingual VSR model on LRS3 and mTEDx test set. The Red and Blue indicate deletion and wrong predicted words, respectively.